Listed are the calories and grams of saturated fat for a sample of ten food items. Find the following a. Mean calories b. Median calories c. Mean grams of saturated fat d. Median grams of saturated fat Grams of Food item Calories saturated fat Chocolate bar (1.45 ounces) Meat & veggie pizza (large slice) New England clam chowder (1 cup) 216 7.0 364 5.6 1.9 149 Baked chicken drumstick 75 0.6 (no skin, medium size) Curly fries, deep-fried (4 ounces) Wheat bagel (large) 3.2 276 375 0.3 Chicken curry (1 cup) 146 1.6 0.5 Cake doughnut hole (one) 59 Rye bread (1 slice) 67 0.2 Raisin bran cereal (1 cup) 0.3 195
For the data from the previous problem, do the following Find the standard deviation of calories b. Interpret the standard deviation for calories c. Find the standard deviation for grams of saturated fat d. Interpret the standard deviation for grams of saturated fat a. Using the data from the previous exercise, calculate the coefficient of variation for: a. Calories b. Grams of saturated fat
Using the calories data do the following a. Calculate the z-score for the chocolate bar b. Interpret the z-score for the chocolate bar Use the IQR method to identify any outliers for grams of saturated fat.
Answer #1

A. Mean calories = sum of calories / 10 = 192.2

B. Arrange calories in ascending order, median = average of 5th and 6th value = (149+195)/2 = 172

C. Mean gram of saturated fat = sum of gram of saturated fat / 10 = 2.12

D. Arrange gram of saturated fat in ascending order, median = average of 5th and 6th value = (0.6+1.6) /2 = 1.1